# The Democratization of Wealth Management: Hedged Mutual Fund Blockchain Protocol

### v1

@@ 4.3 Performance Fees Across Aggregated Positions
    • This happens when when the NAV is below the HWM and transactions - or positions - are made when the weighted average NAV is above the NAV. Then if the NAV goes up - but still remains lower than the weighted average NAV - and some withdraw transactions happen they will not get charged a performance fee. These positions that are getting redeemed have enjoyed positive performance since the price they entered is lower than the price they are exiting, but they are paying a performance fee.
@@ 4.4 Performance Fees Across Separate Positions
    Lower HWM and Compensate Older Positions: If the slump continues for a long time and we do not wish to utilize Option (2). The third option is that we lower the HWM and compensate the older positions - and corresponding transactions across investors - based on their point of entry and the HWM for which they have paid the fees. The way to do this would be to calculate for each transaction its point of entry and the corresponding performance fees. Then some compensation scheme for them has to be worked out. This could be an airdrop, or, any scheme with a reward plan that extends over several months (Harrigan et al., 2018; Li et al., 2024; Allen et al., 2024; End-note 16). This can be very expensive both on the treasury and also in terms of gas fees, manual intervention needed perhaps etc.

### v2

@@ 4.3 Performance Fees Across Aggregated Positions
    • This happens when when the NAV is below the HWM and transactions - or positions - are made when the weighted average NAV is above the NAV. Then if the NAV goes up - but still remains lower than the weighted average NAV - and some withdraw transactions happen they will not get charged a performance fee. These positions, that are getting redeemed, have enjoyed positive performance since the price they entered is lower than the price they are exiting, but they are paying a performance fee.
@@ 4.4 Performance Fees Across Separate Positions
    Lower HWM and Compensate Older Positions: If the slump continues for a long time and we do not wish to utilize Option (2). The third option is that we lower the HWM and compensate the older positions - and corresponding transactions across investors - based on their point of entry and the HWM for which they have paid the fees. The way to do this would be to calculate for each transaction its point of entry and the corresponding performance fees. Then some compensation scheme for them has to be worked out. This could be an airdrop, or, any scheme with a reward plan that extends over several months (Harrigan et al., 2018; Li et al., 2024; Allen et al., 2023; End-note 16). This can be very expensive both on the treasury and also in terms of gas fees, manual intervention needed perhaps etc.